# BooookScore: A systematic exploration of book-length summarization in the era of LLMs

### v2

@@ 6 Limitations
Since computing BooookScore requires iterating through a summary sentence by sentence using GPT-4, it can be expensive and slow especially given that the annotation prompt is long (see Appendix J.4). We did experiment with an approach that asked GPT-4 to annotate errors in the entire summary at once, but the generated annotations would often include too many trivial questions, and alignment with human judgments was low. That said, despite the API costs of GPT-4 and the relatively slow time to evaluate one summary, BooookScore is still significant cheaper and faster than performing human evaluations.
@@ 8 Conclusion
Our work presents the first systematic study of book-length summarization using LLMs. We establish a novel human evaluation protocol to assess summary coherence on newly-published books. Then, we develop an LLM-based automatic metric called BooookScore that relies on a coherence error taxonomy derived from our human annotations. Using BooookScore allows us to evaluate various prompting strategies and model choices, revealing insights such as: hierarchical merging produces more coherent summaries but may lack detail compared to incremental updating; closed-source models like GPT-4 and Claude 2 generate the most coherent summaries; and increasing chunk size can significantly improve incremental updating. Interesting future directions include automatically evaluating faithfulness in the book-length summarization setting, benchmarking newer long-context LLMs using BooookScore, and expanding BooookScore to multilingual texts. We release our BooookScore metric and annotated summaries to enable meaningful progress in book-length summarization.
@@ A.2 Incremental updating
    Iterate through the remaining chunks. If $g_{i}$ exceeds the maximum summary length $G_{n}$, call the compression prompt to compress $g_{i}$ to fit within the length limit. [^20] As the final summary often contains artifacts like “in the current segment” or “in this section”, especially with weaker models, we included an additional post-processing prompt to clean up these artifacts. See Appendix J.3. See Appendix A.2.1 for more details on compression.
@@ D.1 Data processing details
In order to preserve all visual separators within the text, we extract all text elements from the epub files without further automatic processing. As a result, sometimes content from non-narrative sections would appear in the generated summaries. We apply simple post-processing by prompting GPT-4 to remove information coming from non-narrative sections of the book. The prompt can be found in Appendix J.3.
Appendix E Label-wise alignment between GPT-4 and human annotations
@@
Appendix F Effectiveness of existing reference-free evaluation metrics
@@
LLaMA 2 (2048) 0.013 0.3446
We investigate how well existing reference-free evaluation metrics work for our setting where the source text is at book level. BLANC (Vasilyev et al., 2020) measures how helpful a summary is to understanding the source document by testing if a pre-trained model can better fill in masked words given access to the summary. SUPERT (Gao et al., 2020) measures the semantic similarity between the a summary and some pseudo reference summaries generated from the source document. We compute both metrics for hierarchical summaries generated under all configurations, and show results in Table 5. Both BLANC and SUPERT differ by nearly negligible margins across models, which makes them less meaningful. Furthermore, both metrics rate Claude 2 (88000) summaries as inferior to ChatGPT (2048) summaries, a finding that clearly contradicts results from our qualitative analysis.
@@
Appendix G Computing annotation precision
@@
We mix human and GPT-4 annotations in the data we present to the annotators, concealing the origin of these annotations to prevent any potential bias. Each annotator was responsible for the human and GPT-4 annotations of 25 books (different from the 25 books which they annotated summaries for). We paid \$1.7 per summary and \$0.15 per annotation. 100 summaries and 1659 annotations resulted in a total cost of \$418.85.
@@ Appendix H API costs
Total 4024.3 5165
Appendix I Example summaries
@@
Table 9: ChatGPT 2048 summaries of Janika Oz’s A History of Burning.
@@
Appendix J Prompts
J.1 Incremental updating

### v3

@@ 6 Limitations
Since computing BooookScore requires iterating through a summary sentence by sentence using GPT-4, it can be expensive and slow especially given that the annotation prompt is long (see Appendix M.4). We did experiment with an approach that asked GPT-4 to annotate errors in the entire summary at once, but the generated annotations would often include too many trivial questions, and alignment with human judgments was low. That said, despite the API costs of GPT-4 and the relatively slow time to evaluate one summary, BooookScore is still significant cheaper and faster than performing human evaluations.
@@ 8 Conclusion
Our work presents the first systematic study of book-length summarization using LLMs. We establish a novel human evaluation protocol to assess summary coherence on newly-published books. Then, we develop an LLM-based automatic metric called BooookScore that relies on a coherence error taxonomy derived from our human annotations. Using BooookScore allows us to evaluate various prompting strategies and model choices, revealing insights such as: hierarchical merging produces more coherent summaries but may lack detail compared to incremental updating; and increasing chunk size can significantly improve incremental updating. Interesting future directions include automatically evaluating faithfulness in the book-length summarization setting, benchmarking newer long-context LLMs using BooookScore, and expanding BooookScore to multilingual texts. We release our BooookScore metric and annotated summaries to enable meaningful progress in book-length summarization.
@@ A.2 Incremental updating
    Iterate through the remaining chunks. If $g_{i}$ exceeds the maximum summary length $G_{n}$, call the compression prompt to compress $g_{i}$ to fit within the length limit. [^18] As the final summary often contains artifacts like “in the current segment” or “in this section”, especially with weaker models, we included an additional post-processing prompt to clean up these artifacts. See Appendix M.3. See Appendix A.2.1 for more details on compression.
@@ Appendix D More experiment details
We use the LLaMA-2-7B-Instruct checkpoint [^19] https://github.com/togethercomputer/Llama-2-7B-32K-Instruct fine-tuned on long-context summarization (BookSum). For Mixtral, we used the mistralai/Mixtral-8x7B-Instruct-v0.1 checkpoint hosted by the Together API. For closed-source models, we use the gpt-4 2023-03-15 and gpt-3.5-turbo-0301 checkpoints on Microsoft Azure. Anthropic unfortunately does not disclose checkpoint information, but our summaries were all obtained via their Claude 2 API in September 2023.
@@ D.1 Data processing details
In order to preserve all visual separators within the text, we extract all text elements from the epub files without further automatic processing. As a result, sometimes content from non-narrative sections would appear in the generated summaries. We apply simple post-processing by prompting GPT-4 to remove information coming from non-narrative sections of the book. The prompt can be found in Appendix M.3.
Appendix E Experiments with SQuALITY
To investigate the effect of incremental updating and hierarchical merging on summary coherence, we evaluated GPT-4 on the validation set of the SQuALITY dataset, which contains sci-fi stories that are 4000-6000 words long. Table 5 shows the ROUGE-L scores. The baseline setting is where the model summarizes the stories in one go, truncating the stories whenever they exceed the model’s context window. Using incremental updating lowers ROUGE-L by a small amount, indicating that baseline summaries have slightly more overlap with the provided reference summaries. We will update the BooookScore of these summaries in the next version of the paper.
Table 5: ROUGE-L of summaries generated by GPT-4 on the SQuALITY validation set under two settings: baseline and incremental updating.
Chunk size ROUGE-L
Baseline
- 14.9
Incremental updating
256 13.2
512 13.7
E.1 Qualitative analysis
We show the baseline, incremental, and human summaries for Venus Is a Man’s World from SQuALITY in Table 6. The incremental summary was generated by GPT-4 with a chunk size of 512.
The baseline summary provides the most coherent and comprehensive overview of the story’s key events and themes. It clearly introduces the main characters - Ford, Evelyn, and Butt - and their backgrounds, while succinctly summarizing the plot including Ford’s friendship with Butt, Evelyn’s discovery and confrontation, and Butt’s unconventional marriage proposal. The summary highlights central conflicts related to gender roles and norms. In contrast, the incremental summary, while touching on similar plot points and themes, does so in a more disjointed, less cohesive manner. Details about the setting, dunging drug harvesting, and weaponry feel redundant. Finally, the human summary lacks clarity in many parts, with confusing references to the Male Desuffrage Act and inconsistent character details. It focuses heavily on the early parts of the plot at the expense of later key events.
Overall, our finding that GPT-4 summaries can be better than human summaries aligns with results from recent work on LLM-based short-document summarization (Pu et al., 2023b). The qualitative analysis shows that methods like incremental updating could introduce coherence errors into LLM-generated summaries.
Table 6: Summaries of William Tenn’s Venus Is a Man’s World.
Baseline Incremental updating Human
Venus Is a Man’s World is a story about a young boy named Ferdinand, or “Ford,” and his sister Evelyn, who are traveling on a spaceship called the Eleanor Roosevelt to the planet Venus. Evelyn is seeking a husband on Venus, as there is a shortage of men on Earth due to wars and other factors. Ford, curious and adventurous, explores the ship and stumbles upon a stowaway named Butt Lee Brown, a Venusian farmer who had gotten into trouble on Earth and is now hiding in a lifeboat. Ford and Butt form a bond, with Butt teaching Ford about Venusian culture, songs, and even politics. Ford starts to admire Butt’s strong, independent nature and begins to question the Earth’s female-dominated society. However, when Evelyn discovers Ford’s friendship with Butt, she confronts the stowaway, accusing him of corrupting her brother’s morals and breaking the law by stowing away on the ship. Butt, unfazed by Evelyn’s accusations, proposes marriage to her, stating that they are both on the same journey to find a partner and that she has the qualities he is looking for in a wife. Ford excitedly urges his sister to accept Butt’s proposal, but the story leaves the outcome of this unconventional proposal unresolved. The story highlights the clash of gender roles, societal norms, and the challenges faced by individuals seeking love and companionship in a rapidly changing world. In the story, siblings Ford and Evelyn from an underwater city accompany women seeking husbands on a journey to Venus. Ford befriends stowaway, Butt Lee Brown, who has a troubled past. They bond over their life stories as Butt educates Ford about Venusian culture and weaponry. Despite being outsiders, they travel together. Evelyn convinces Ford to attend a geography lecture, where she becomes deeply engaged in the discussion of illegal dunging drug harvesting on the Macro Continent. Ford’s knowledge of dunging and the Venusian term “Paddlefoot” piques Evelyn’s interest. Ford argues that Venusians are pioneers who need strong, independent men for building a civilization on the new planet, thus intriguing Evelyn. Ford takes Evelyn to meet Butt Lee Brown, making her promise not to turn him in. Upon meeting, Evelyn confronts Butt about his illegal activities and moral delinquency. Butt dismisses her concerns, sparking a debate between them about Earth and Venusian societies. During their conversation, Butt proposes to Evelyn, stating that they both have similar goals in finding a spouse and that she possesses qualities he finds desirable. Evelyn is taken aback by the proposal, and her brother Ford excitedly encourages her to accept. The story continues to explore the characters’ relationships and their differing perspectives on the new planet, as well as the potential consequences of their actions and decisions. Ferdinand is on a spaceship with his elder sister. The ship is jam-packed with females going to Venus in search of husband and he exits the cabin while the women are still in their hammocks. The women at his times have all the rights and do all the important things since the Male Desuffrage Act, so the boy is admitted to the ship on behalf of his sister. He explores the empty ship in search of portholes and after some hesitation enters a forbidden area. There he looks at the stars and then tries to open a lock on the lifeboat. A huge scary man appears with a blaster and frighteningly cold gaze. Ferdinand explains that he comes from Undersea, an area on Earth, and tells his family story - his parents being one of the first married couples in Undersea and dying a while ago, leading to his sister’s decision to migrate to Venus. The stranger, Butt, tells about the lack of women on Venus and his travel to Earth in search of a wife without any idea “it’s a woman’s world”. So he got in trouble with the law and stowed away on this ship. His many brothers were killed in a rising and only one is left. From that day on Ferdinand keeps visiting the stowaway bringing fruits and listening to stories about Venus. Butt teaches the boy to use the blaster without giving it not hold and constantly asks about Evelyn, the sister. Once, Ferdinand attends a geography lecture on the ship with his sister and corrects the lecturer about Venusian geography. Evelyn starts eliciting where the boy learned that and the boy tells about real men working on Venus. Sis gets angry with those masculine ideas and doesn’t believe them to come from a little boy. Ferdinand tries to lie but Sis suppresses him into confession and he leads her to Butt. She tells Butt about all the laws he has broken while the least responds with an appeal to sense. Suddenly, Butt simply proposes a mutually beneficial marriage to stop the debate.
Appendix F Effect of summary length on BooookScore
Figure 3: BooookScore vs. summary length.
To explore the impact of summary length on BooookScore, we plot BooookScore vs. length in Figure 3 for the 100 summaries generated by GPT-4 with incremental updating using a chunk size of 2048. The plot indicates that there is no discernible correlation between length and BooookScore.
Appendix G Bootstrapping analysis of BooookScore
To check the stability of the BooookScore metric, we ran a bootstrapping experiment. Given the BooookScore of 100 summaries generated by GPT-4 using incremental updating with a chunk size of 2048, we randomly sample 1000 times with replacement using a sample size of 100, then compute the mean of these 1000 samples. The standard deviation of these samples is 0.015, which suggests that BooookScore is consistent and reliable across multiple random samples.
Appendix H Label-wise alignment between GPT-4 and human annotations
@@
Appendix I Effectiveness of existing reference-free evaluation metrics
@@
Mixtral-8x7B (2048) 0.0169 0.3836
LLaMA2-7B-Inst (2048) 0.013 0.3446
We investigate how well existing reference-free evaluation metrics work for our setting where the source text is at book level. BLANC (Vasilyev et al., 2020) measures how helpful a summary is to understanding the source document by testing if a pre-trained model can better fill in masked words given access to the summary. SUPERT (Gao et al., 2020) measures the semantic similarity between the a summary and some pseudo reference summaries generated from the source document. We compute both metrics for hierarchical summaries generated under all configurations, and show results in Table 7. Both BLANC and SUPERT differ by nearly negligible margins across models, which makes them less meaningful. Furthermore, both metrics rate Claude 2 (88000) summaries as inferior to GPT-3.5-Turbo (2048) summaries, a finding that clearly contradicts results from our qualitative analysis.
@@
Appendix J Computing annotation precision
@@
We mix human and GPT-4 annotations in the data we present to the annotators, concealing the origin of these annotations to prevent any potential bias. Each annotator was responsible for the human and GPT-4 annotations of 25 books (different from the 25 books which they annotated summaries for). We paid \$1.7 per summary and \$0.15 per annotation. 100 summaries and 1659 annotations resulted in a total cost of \$418.85 (USD). Note that annotators were informed of this evaluation task weeks after they completed the first task (where they were asked to highlight spans and ask questions given book summaries). Thus, monetary reward for the second task could not have biased annotators to produce as many annotations as possible for the first task.
@@
Mixtral-8x7B 2048 - 395.7
Total 4020.5 6021.2
Appendix L Example summaries
@@
Table 11: GPT-3.5-Turbo 2048 summaries of Janika Oz’s A History of Burning.
@@
Table 14: Mixtral 2048 summaries of Janika Oz’s A History of Burning.
Strategy Summary
Hier ”A History of Burning” by Janika Oza is an intricate, multigenerational saga that explores themes of intergenerational trauma, resilience, identity, and colonial exploitation across time, spaces, and continents. The narrative spans from Jamnagar, Gujarat, to Kenya, Uganda, and Canada, weaving together the lives of several characters whose experiences intersect and diverge amidst political upheaval, personal sacrifices, and cultural expectations. The story begins in Jamnagar during a severe drought, where thirteen-year-old Pirbhai struggles to find work. Driven by his mother’s encouragement, he embarks on a journey to Jamnagar and meets a mysterious merchant who leads him to work on a dhow sailing towards an unknown destination. Meanwhile, in Kenya, Sonal is a young Indian woman living with her family. When Pirbhai, now a refugee in Kenya, is employed by Sonal’s father, the two form a deep bond as they navigate their lives together amidst political turmoil and personal sacrifices. Pirbhai’s past unveils his harrowing experiences as an indentured laborer in East African colonies and the guilt he feels over the circumstances that led him to Kenya. Simultaneously, the narrative explores Sonal’s losses, including her brothers who perished during migration, and her struggle to forge connections with her family and her new home. The narrative intertwines with Rajni’s story during her migration from Uganda to Canada, following Idi Amin’s expulsion of Asians from the country. Rajni, a resilient woman, grapples with her new life in a cold and foreign country while mourning the loss of her daughter, Latika. The lives of these characters continue to intersect and diverge in Toronto, revealing a rich tapestry of experiences, hopes, and triumphs despite the adversities they encounter. Other characters play significant roles in shaping the narrative. Vinod, an Indian immigrant, navigates cultural assimilation, identity, and family expectations in Canada. Kiya, Rajni’s daughter, deals with her resentment towards her sister Mayuri, who left the family during their time of need. Hari, the Patels’ son, and his friend Solomon grapple with the consequences of their actions during a protest against racism. Through the characters’ experiences, the narrative explores the intricate connections between historical events and their impact on the characters’ lives. From the struggles of poverty in Jamnagar, the brutal labor conditions in East African colonies, to the racial tensions and political turmoil in Uganda, the story emphasizes the power of resilience, human connections, and the complex nature of identity forged amidst shifting political alliances and dislocations. In Toronto, Kiya joins the army, and the family grieves after Latika’s disappearance. Hari and Solomon confront the consequences of their actions during a protest against racism. The non-linear storytelling features past and present flashbacks, providing a rich context for the characters’ experiences. Despite the non-linear narrative, the story ultimately presents a consistent and chronological account of the characters’ lives and their struggles to forge a sense of belonging amidst adversity. In conclusion, ”A History of Burning” is a compelling exploration of characters’ lives woven together throughout history, emphasizing the profound power of resilience, human connections, and the complex nature of identity. By engaging with themes such as colonial exploitation, community, survival, and identity, the story forges a rich and imaginative world that touches on the lives of Pirbhai, Sonal, Rajni, Vinod, Kiya, Mayuri, and others, as they navigate across continents and navigate an intricate web of cultural expectations, personal sacrifices, and political upheaval.
Inc The story follows the Pirbhai family, whose roots lie in India but have ties to Kenya, Uganda, and now Canada. The family includes Rajni and Vinod, their sons Hari and the late Arun, daughter Mayuri, Rajni’s uncle Nanu, reemerged sister Kiya, and her son Harilal. In 1991, Hari bonds with friends Marcus, Solomon, and Sami over shared experiences and basketball. He grapples with Arun’s death and the revelation of a long-lost sister, Latika. The family unites to discuss Latika and commemorate Arun. Hari investigates the Ugandan Asian expulsion but discovers limited information. In 1992, Hari turns 19 and forms a close friendship with Marcus, Solomon, and Sami. Ignited by the Rodney King beating and the deaths of Raymond Lawrence and Michael Wade Lawson, he joins community protests, driven by his pursuit of justice and skepticism of official narratives. Together, they restore a vandalized local basketball court. Hari’s reflections on family dynamics lead him to play basketball with Solomon instead of completing a report. Parallelly, Rajni navigates Vinod’s deteriorating health while embracing his desire for privacy. Together, they face his physical decline. Rajni’s past reemerges while she raises her concerns for Hari’s safety during protests. She understands the importance of supporting their causes despite her fears and apprehensions. At the demonstration, Hari advocates for justice and unity. When tensions escalate, he gets separated from his friends. Rajni turns to prayer for Hari’s safety, prompted by reminiscences of her past. Upon discovering Hari’s disappearance, the family unites in searching for him. After the protest, disoriented and lost, Hari reconnects with his friends, who assist him in returning home. Aware of Hari’s safety, the family reconvenes to make sense of the day’s events, solidifying their unity and commitment to fighting for justice. The narrative intertwines Rajni, Hari, and their friendships as they confront their histories, find comfort in unity, and sustain their roots amid political strife. Acclaimed writers like Sharon Bala, Ngig) wa Thiong’o, and Gaiutra Bahadur enrich the storyline with dedicated support, historical context, and scholarship. Mentors and writing programs bolstered Hari’s belief in his writing abilities. The Pirbhai family’s story spans generations and continents, with each member experiencing their own trials and triumphs. They carry the weight of their family legacy and the history of their ancestors. Through it all, the family remains united, finding strength in their shared experiences and their commitment to fighting for justice and equality. The story emphasizes the importance of community, collectivism, and solidarity, inviting readers to reflect on who is included and excluded in their own spheres.
@@
Appendix M Prompts
M.1 Incremental updating